A teacher has two students, one two times as old as the other. The teacher is five times as old as the older student, and in 5 years she will be five times as old as her younger student. Find the teacher’s present age.
Answer by Edwin McCravy(20060)

A teacher

teacher's age = t

has two students, one two times as old as the other.
younger student's age = s
older student's age = 2s 

The teacher is five times as old as the older student,
t = 5(2s)
t = 10s

and in 5 years
teacher's age in 5 years = t+5
younger student's age in 5 years = s+5

she will be five times as old as her younger student.
t+5 = 5(s+5)

Using t = 10s, substitute 10s for t in

t+5 = 5(s+5)
10s+5 = 5(s+5)

Solve that for s, the age of the younger student.
Then substitute what you get for s in

t = 10s

That will be the teacher’s present age.

Edwin